What rate of interest is required to triple an investment in 12 years if the interest is compounded continuously? Use A+Pe^rt.
Answer by lwsshak3(11628)

What rate of interest is required to triple an investment in 12 years if the interest is compounded continuously? Use A+Pe^rt.
***
A=Pe^rt, P=initial investment, r=rate of interest, A=amt after t-years
A/P=3=e^rt
rtlne=ln3
lne=1
rt=ln3
r=ln3/12≈0.09155
rate of interest required=9.2% (rounded)